OpenStudy (anonymous):

Complete the square: y=5x^2-4x+2

OpenStudy (amistre64):

y=5x^2 -4x +2 y-2 = 5x^2 -4x K+(y-2)/5 = x^2 -(4/5)x + K where K= [(-4/5)/2]^2 = (-4/10)^2 = 16/100 = 4/25 (4/25) +(y-2)/5 = x^2 -(4/5)x +(16/100) (4/5 +y -2)/5 = (x-(4/10))^2 [y-(6/5)]/5 = (x-(2/5))^2 y-(6/5)= 5(5x-2)^2 y= [5(5x-2)^2] + (6/5) Thats the best I can do with it :)

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

Id get a second opinion just to be safe :)